OpenStudy (anonymous):

A trapezoid is formed by lines with the following equations. 2x+y= 16 x= 4 x= 0 y= 0 Find the area of the trapezoid.

OpenStudy (aum):



OpenStudy (aum):

OABC is the trapezoid. Area of trapezoid = 1/2 * (sum of the parallel sides) * distance between the parallel sides. Area = 1/2 * (OC + AB) * OA OA = 4 because point A is (4,0) and point O is (0,0) Area = 1/2 * (OC + AB) * OA = 1/2 * (OC + AB) * 4 = 2 * (OC + AB) To find OC, we need the coordinates of point C. Point C is the y-intercept of the line 2x+y=16. What is the y-intercept? Point B is the intersection of the line x = 4 and 2x+y=16. Find point B.